OpenStudy (anonymous):

The points (-4,-3) and (-1,-8) are on a line. Find intercepts to the nearest tenth. How would i set up this problem to solve to get one of these answers? A) x=-5.8 y=-9.7 B) x=-5 y=-11 C) x=5.8 y=-9.7 if you show me how to solve this problem.. you'll receive a medal & a fan!

OpenStudy (anonymous):



OpenStudy (anonymous):

that is the equation plug in the point

OpenStudy (anonymous):

points*

OpenStudy (dmndlife24):

Use slope formula m = y2 - y1/x2 - x1 where m = slope Then when u find the slope plug it in the point slope formula y - y1 = m (x - x1) y1 = -3 and x1 = -4 Simplify your answer to standard form and, first, set x equal to 0 to find the y - int. Then, set y equal to 0 to find the x - int.

OpenStudy (dmndlife24):

lets start with finding the slope... y2 = -8 y1 = -3 x2 = -1 x1 = -4 m = (-8) - (-3)/(-1) - (-4) m = -5/3

OpenStudy (anonymous):

i got that part

OpenStudy (anonymous):

but finding the other part, I dont get

OpenStudy (dmndlife24):

Then take that -5/3 and plug it in the point slope formula... y1 = -3 x1 = -4 y - (-3) = -5/3 (x - (-4)) Do you get how i plugged it in @Bri37?

OpenStudy (anonymous):

i get that part & whats the next . i dont understand or know how to do the next part

OpenStudy (dmndlife24):

Okay...next, we simplify what we plugged in... y - (-3) = -5/3 (x - (-4)) y + 3 = -5/3 (x + 4) Do you understand how i got this?

OpenStudy (anonymous):

Got that. i dont understand about how you'll get the right answer

OpenStudy (dmndlife24):

Well, what we have to do next is simplify further and put it in standard form...we do this so we can find the x and y intercepts...